Let f be a continuous function on the interval [a,b], and let c = a+b 2 . Suppose that F is any antiderivative of f with F(a) = 3 and F(b) = 7.
Find: (a)Rb a f(x)dx (2 marks) (b) −Ra b f(x)dx (2 marks) (c)Rc a f(x)dx +Rb c f(x)dx (2 marks) (d) Dx[−(Rc a f(x)dx +Rb c f(x)dx)] (2 marks) 3. Find the area bounded by the curve y = |x−2|, the x-axis and the lines x = −7 and x = 11. (5 marks)
1
Expert's answer
2020-06-10T19:41:42-0400

Let "f" be a continuous function on the interval "[a,b]," and let "c=\dfrac{a+b}{2}." Suppose that "F" is any antiderivative of "f" with "F(a)=3" and "F(b)=7."

Find:


"\displaystyle\int_{a}^bf(x)dx"

By the Fundamental Theorem of Calculus (Part 2) if "f"  is continuous on "[a, b]"  then


"\displaystyle\int_{a}^bf(x)dx=F(b)-F(a),"

where "F" is any antiderivative of "f."

Given "F(a)=3" and "F(b)=7." Then


"\displaystyle\int_{a}^bf(x)dx=F(b)-F(a)=7-3=4"

b)


"-\displaystyle\int_{b}^af(x)dx=-(-\displaystyle\int_{a}^bf(x)dx)=\displaystyle\int_{a}^bf(x)dx=4"

c)


"\displaystyle\int_{a}^cf(x)dx+\displaystyle\int_{c}^bf(x)dx=\displaystyle\int_{a}^bf(x)dx=4"

d)


"D_x[-(\displaystyle\int_{a}^cf(x)dx+\displaystyle\int_{c}^bf(x)dx)]=D_x[-4]=0"

2.


"y=|x-2|= \begin{cases}
 x-2 &\text{if } x\geq 2 \\
 -x+2 &\text{if } x<2
\end{cases}"

"Area=A=\displaystyle\int_{-7}^2(-x+2)dx+\displaystyle\int_{2}^{11}(x-2)f(x)dx="

"=[-{x^2\over 2}+2x]\begin{matrix}
 2 \\
-7
\end{matrix}+[{x^2\over 2}-2x]\begin{matrix}
 11 \\
2
\end{matrix}="

"=-2+4+{49\over 2}+14+{121\over 2}-22-2+4=81(units^2)"

Area is "81" square units.